# Let the Trial Begin: A Mock-Court Approach to Vulnerability Detection using LLM-Based Agents

### v1

@@ 8. Conclusion and Future Direction
In this study, we proposed VulTrial, a novel multi-agent LLM-based approach for vulnerability detection that applies a courtroom-style paradigm. VulTrial comprises of four specific-role agents with distinct roles: the security researcher (prosecutor), code author (defense attorney), moderator (judge), and review board (jury). Each agent is essential to get the best performance (Section 4.4). We evaluated VulTrial against single-agent (Ding et al.’s CoT) (Ding et al., 2025b) and multi-agent, GPTLens (Hu et al., 2023), vulnerability detection baselines. Our results show that VulTrial outperforms both the single-agent and multi-agent baselines. In experiments using GPT-3.5, VulTrial enhances the performance of the single-agent baseline by 151.69%, and outperforms GPTLens by 239.78%. Additionally, instruction tuning with just 50 sample pairs yields further performance improvement. When utilizing GPT-4o with instruction tuning in moderator agent, VulTrial improves single-agent and multi-agent baselines by 139.89% and 118.3%. Furthermore, the number of discussion rounds among the first three agents in VulTrial does not improve the overall performance.
@@
For future direction, we plan to apply VulTrial to different domains to evaluate its generalizability further. We also plan to test additional LLMs such as DeepSeek-R1 (DeepSeek-AI et al., 2025) and LLaMA3 (Grattafiori et al., 2024) and consider each model’s strengths and weaknesses to choose the one that best fits for each role-specific task.

### v2

@@ 8. Conclusion and Future Work
We proposed VulTrial, a novel explainable multi-agent LLM-based approach for vulnerability detection that applies a courtroom-style paradigm. VulTrial comprises four role-specific agents: the security researcher (prosecutor), code author (defense attorney), moderator (judge), and review board (jury). Each agent is essential to get the best performance (Section 4.4). We evaluated VulTrial against single-agent (CodeBERT (feng2020codebert), CodeT5 (wang2021codet5), UniXCoder (guo2022unixcoder), LineVul (fu2022linevul), and Ding et al.’s CoT (ding2024vulnerability)) and multi-agent (GPTLens (hu2023large)) baselines. Our results show that VulTrial outperforms both the single-agent and multi-agent baselines, with instruction tuning in the moderator agent further enhancing its performance. We have also demonstrated the efficacy of VulTrial across different LLMs, including an open-source, in-house-deployable model (LLaMA-3.1-8B), as well as the high quality of its generated explanations and its ability to uncover multiple zero-day vulnerabilities.
@@
For future work, we plan to apply VulTrial to more datasets to evaluate its generalizability further. We also plan to consider the strengths and weaknesses of different LLMs to choose the one that best fits as the base model for each role-specific task.
Acknowledgements.
This research / project is supported by the National Research Foundation, Singapore, and the Smart Nation Group under the Smart Nation Group’s Translational R&D Grant (Award No. TRANS2023-TGC02). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore or the Smart Nation Group.